# Gamma-Ray Burst Light Curve Reconstruction: A Comparative Machine and Deep Learning Analysis

### v4

@@ 2.3 The Machine-Learning Approach
Machine learning models are algorithms that learn from the data, recognize patterns, and give predictions on unseen data. One of the most widely used ML techniques is the Artificial Neural Network (ANN) (Heaton, 2018; Wang, 2003). These are inspired by how the human brain (neurons) processes information. Just as the brain consists of interconnected neurons that help us think and learn, a neural network consists of layers of artificial neurons that process and learn from data. A neural network consists of the input, hidden, and output layers. The input layer is the first layer that receives the raw input data. In this layer, no computations are performed, as it is responsible for passing the data to the next layer. The output layer is the final layer that gives the output predictions of the hidden layer. The hidden layers consist of artificial neurons to analyze the data. Each connection between neurons has a weight, determining the input value’s importance. A bias is an extra value added to fine-tune and adjust the predictions. At first, the weights and biases are initialized randomly, but they are adjusted to improve the prediction as the model trains. Training happens in multiple rounds (epochs), and after each epoch, the model refines the weights and biases depending on the errors it makes.
@@
    This resulted in a fixed set of hyperparameters (including the number of epochs) for each model architecture. These were then used to train each model on the complete dataset of 545 GRBs. This tuning process was performed separately for each model type evaluated in the study.
@@
    • Attention UNet: Number of filters per layer, kernel size, number of convolutional blocks, learning rate, batch size, number of epochs
@@
  • Min-Max Scaling: For each GRB, the data is scaled to the [0, 1] range using standard min-max normalization (4). However, scaling is not applied to UNetAttention and MLP models, as by performing some tuning in several GRBs showed a decrease in performance. Additionally, UNetAttention inherently includes multiple Batch Normalization layers, while the MLP, being a relatively simpler network, is less susceptible to the weights of the model either increasing exponentially or going toward zero, namely exploding or vanishing gradients.
@@
  • 5-Fold Cross-Validation: After fixing the hyperparameters for each model, 5-fold CV is used as an alternate metric to evaluate the performance of each model. Rather than splitting the entire dataset of 545 GRBs into separate training and testing sets, we apply CV to each GRB separately. For each GRB, its time-flux data points are first randomly shuffled and then partitioned into five approximately equal-sized, non-overlapping subsets (folds). In each of the five iterations, four folds (80% of the data) are used to train the model, and the remaining fold (20%) is used to validate it (Kohavi, 1995). This ensures that all data points are used for both training and validation across the five folds. After training, the model predicts flux values at the time points in the validation fold. These predicted fluxes are inversely transformed back to the original scale before calculating the MSE, as it depends on the scale. The MSE loss is averaged across the folds and then across the 545 GRBs. CV was used solely to evaluate model performance, not for tuning model parameters or minimizing training error.
@@ 2.3.1 Gaussian Process
GP regression (Rasmussen, 2003) is a probabilistic, non-parametric approach designed for regression tasks. The probabilistic nature of the GP model allows for the likelihood of the prediction instead of producing a singular prediction. The method relies on Bayesian inference, where the prior based on trends in the observed data is updated to compute a posterior that aligns with the data. The model prediction is bound by the confidence interval, which defines the region of likelihood of the prediction.
@@
In our data, out of the total sample of 545 GRBs, a tiny subset containing 9 GRBs with redshifts and 15 GRBs without redshifts, exhibited significant temporal gaps, defined as time intervals greater than 0.5 (in log10 scale). As Fig. 2a shows, the GP Regressor overfitted these data points in both the gaps, resulting in erratic predictions and extensive 95% confidence intervals. This inconsistent behavior suggested the need to fine-tune two primary hyperparameters: alpha and n_restarts_optimizer, by employing GridSearchCV; alpha is a value added to the kernel matrix’s diagonal during fitting. This hyperparameter helps avoid potential numerical issues, shown in Fig. 2a, by ensuring the matrix remains positive and definite. Furthermore, alpha can be interpreted as the additional Gaussian measurement noise variance on the training observations. The default value is 1e-10. A more considerable alpha value is frequently required to avoid overfitting in cases where the data is noisy. n_restarts_optimizer denotes the number of times the optimizer restarts to find the kernel parameters that maximize the log-marginal likelihood. This procedure guarantees a more reliable optimization of the kernel’s parameters and aids in avoiding local minima.
@@
After tuning the hyperparameters, we fit the GP Regressor to the data. The optimized model improved the reconstruction of the LCs, as shown in Fig. 2b. It might appear that Fig. 2a offers a better reconstruction than Fig. 2b, as it reconstructs the flare in the first temporal gap. However, this is misleading as very large flares like this have been unusual in the sample of the 545 GRBs. Additionally, the confidence intervals in Fig. 2a are significantly broader due to the larger temporal gap, reflecting high uncertainty. After hyperparameter optimization, the model shown in Fig. 2b not only yields smoother and more stable predictions but also exhibits reduced uncertainty, as reflected in narrower confidence intervals.
@@ 2.3.2 Multi-layer Perceptron Model
The kernel initialization of each hidden layer uses the He-Uniform (He et al., 2015) initializer method. The He-Uniform initializer initializes weights by sampling from a uniform distribution within the range:
@@
here, $n_{in}$ is the number of input units (neurons) to the layer. This ensures that the weights are small enough to prevent exploding gradients while still being large enough to avoid vanishing gradients. It is particularly suited for ReLU activation function (He et al., 2015) since it does not saturate like sigmoid or tanh activation functions. The ReLU activation function, defined as $max(0,x)$, outputs zero for negative values and keeps positive values unchanged, making it efficient for training deep networks.
@@
The model is trained using the Adam optimizer (Kingma & Ba, 2014) with 0.001 as the learning rate and MSE as the loss function. The number of epochs is set to 5000. The model was implemented using the TensorFlow framework with Keras.
Next, we perform the 5-fold CV for all 545 GRBs individually and then report the average Mean Squared Error (MSE) across all GRBs as our performance metric in Table 5.
@@ 2.3.3 Attention U-Net
U-Net is a convolutional neural network architecture designed primarily for biomedical image segmentation (Ronneberger et al., 2015) but has been successfully adapted for one-dimensional data such as time series (Perslev et al., 2019). The architecture consists of a symmetric encoder-decoder structure, where the encoder compresses the input into a latent representation, and the decoder reconstructs the output from this latent space. Latent space is a lower-dimensional representation of the data where only the key features are retained and the rest discarded. Skip connections between the encoder and decoder ensure the preservation of spatial details, which are critical for high-resolution reconstruction tasks.
For the LCR, we enhance the standard U-Net with an attention mechanism(Oktay et al., 2018), which is a section of a neural network that enhances feature extraction and boosts performance on tasks involving complicated or sequential data by dynamically focusing on the most pertinent portions of the input data and giving distinct aspects varied degrees of priority as stated in Vaswani et al. (2017). The attention mechanism enables the model to focus on relevant features at each resolution level by weighting them based on their importance. This allows the network to capture both global and local structures effectively.
@@
Attention $\text{Attention Output}=x\cdot\psi_{f}.$ (15)
@@
The decoder reconstructs the temporal sequence by progressively restoring the temporal resolution through upsampling layers, which increases the resolution of the input data. Skip connections between the encoder and decoder layers ensure that fine-grained temporal details are preserved. To enhance the utility of skip connections, attention blocks are applied (Oktay et al., 2018).
@@
The U-Net with Attention mechanism is well-suited for tasks involving complex temporal structures like GRB LCs, where capturing long-range dependencies and local variations is critical (Perslev et al., 2019).
@@ 2.3.4 Long Short Term Memory Neural Network
Long Short-Term Memory (LSTM; Hochreiter & Schmidhuber 1997) is a form of Recurrent neural network (Medsker et al., 2001) that performs well in sequence prediction through its ability to capture long-term dependencies. Unlike traditional neural networks, LSTM incorporates feedback connections, allowing it to process data sequences rather than individual data points. This distinctive architecture equips LSTM with the ability to effectively learn and retain crucial information over time, making it a competent deep learning tool well-suited for applications in speech recognition, language translation, time series forecasting, and more (Zhang et al., 2016; Cheng et al., 2016; Hua et al., 2019).
@@
LSTM networks have proven valuable in astrophysics, particularly for time-series data analysis of dynamic celestial phenomena. An example is the LSTM-FCNN model, which combines LSTM and a fully connected neural network (FCNN) to measure time delays in strongly lensed Type Ia supernovae, crucial for measuring the Hubble constant (Huber & Suyu, 2024). Furthermore, LSTMs have been applied for solar flare prediction from active regions (Liu et al., 2019), effectively predicting them using time-series analysis. These examples show that LSTMs can outperform traditional methods in certain astrophysical time-series tasks.
Our model uses Bi-LSTM (Graves et al., 2005) to reconstruct the GRB LCs. The Bi-LSTM architecture extends the standard LSTM by processing the input data sequence in both forward and reverse directions, allowing it to capture temporal dependencies more effectively. This is particularly useful for GRB LCs, which feature complex temporal patterns such as flares and breaks. Bi-LSTMs also excel at handling missing data by learning from surrounding observations, enabling accurate imputation. The basic structure of the LSTM is described below and shown in Fig.3c. Table 2 defines all the variables and symbols.
@@ 2.3.5 Bi-Mamba Model
Mamba (Gu & Dao, 2024) is a novel sequence modeling architecture based on state space models (SSMs), designed to efficiently capture long-range dependencies while reducing computational complexity. Unlike long short-term memory networks (LSTMs), which rely on explicit gating mechanisms to regulate memory updates, Mamba employs a structured state space framework that enables selective long-range memory retention. This structured recurrence mechanism allows Mamba to model temporal dependencies effectively, achieving linear time complexity, making it particularly suitable for large-scale astronomical datasets such as GRB LCs, which exhibit intricate temporal variations.
@@
Since GRB LCs are discrete time-series data, the Mamba is adapted to such inputs using a Zero-Order Hold (ZOH) method – as described in (Pechlivanidou & Karampetakis, 2022) and (Gu & Dao, 2024) – which converts discrete inputs into piecewise constant signals. The resulting discrete-time formulation is given by:
@@
We used a modified version of the MambaModel from Thielmann et al. (2024) (shown in Fig. 3d), which contains an output activation layer with LeakyReLU activation function and a linear mapping at the end. The model was built using the PyTorch framework with hyperparameter tuning using Optuna. The data (D) was scaled to a range of [0, 1] using MinMaxScaler, given in the Eq. 4. The hidden state dimension (d_state) was set to 32, while the hidden layer expansion factor (d_expand) was configured to 80. A dropout rate of 0.1 was applied to mitigate overfitting. Training was conducted over 120 epochs with a batch size of 64, using the Adam optimizer with a learning rate of 1e-4.
@@ 2.3.7 Conditional Generative Adversarial Networks
GANs Goodfellow et al. (2014) provide an innovative approach to reconstruct the LCs. Unlike traditional models that rely on Markov chains—where the current state is dependent on the previous one—GANs bypass this sequential dependency during the learning phase (Mirza & Osindero, 2014), incorporating it only during gradient backpropagation. These networks are referred to as generative models and are particularly useful when dealing with complex data distributions, reducing overfitting by incorporating a dual-model approach.
@@
The CGANs Mirza & Osindero (2014) extend this architecture by incorporating a condition set $y$ into the generator and discriminator, guiding the data generation process. The architecture of CGAN (as seen in Fig. 3f) comprises a generator $G(z|y)$ with two inputs: a condition set $y$ and a random noise vector $z$ and a conditioned discriminator $D(x|y)$ on $y$. The generator aims to create new samples that closely resemble the original samples by using the noise and the condition set as input. After the generation of samples, these, along with the real samples, are fed into the discriminator. The objective for the discriminator is to effectively classify between the generated and the real samples, retrospectively training the generator and discriminator in the process. It is mathematically represented as:
@@
In our implementation, the generator begins with a dense layer containing 256 neurons and the Leaky-ReLU activation function (Xu et al., 2015). It includes a positive slope $m$ and comparatively a tiny negative slope $\alpha m$, preventing the occurrences of dead neurons. In our case, $\alpha$ is taken to be $21e-$ 2. Following the dense layer, we utilized two Convolution-1D layers with 64 and 16 filter sizes, respectively. Paired with ReLU activation, these Convolution-1D layers enhance the feature extraction process, ensuring distinct, high-quality features. The flux and time values are preprocessed by converting them to a symmetric logarithmic scale. The time values act as condition sets for the CGAN, the flux values provide the actual data, and the noise set contains random uniform noise samples from 0 to 1 covering the whole domain. All layers in the discriminator, except the final binary classification layer, mirror those of the generator. This design ensures a balanced model complexity, preventing either network from overpowering the other. The discriminator loss is computed as the sum of binary cross-entropy losses for correctly classifying real samples as real and generated samples as fake. The generator loss is defined as the binary cross-entropy between the discriminator’s predictions on generated samples and the target label for real samples, thereby encouraging the generator to produce more realistic outputs. This results in reliable data-augmentation capabilities for the reconstruction of time series data (Fu et al., 2019; Brophy et al., 2023).
The proposed method uses TensorFlow and Keras libraries for its implementation, with 5000 epochs and a batch size of 256. The Adam optimizer was used to train the model, optimizing the loss function at a learning rate of $41e-$ 4.

### v5

@@ 2.3 The Machine-Learning Approach
Machine learning models are algorithms that learn from the data, recognize patterns, and give predictions on unseen data. One of the most widely used ML techniques is the Artificial Neural Network (ANN) (DeepLearning; ANN). These are inspired by how the human brain (neurons) processes information. Just as the brain consists of interconnected neurons that help us think and learn, a neural network consists of layers of artificial neurons that process and learn from data. A neural network consists of the input, hidden, and output layers. The input layer is the first layer that receives the raw input data. In this layer, no computations are performed, as it is responsible for passing the data to the next layer. The output layer is the final layer that gives the output predictions of the hidden layer. The hidden layers consist of artificial neurons to analyze the data. Each connection between neurons has a weight, determining the input value’s importance. A bias is an extra value added to fine-tune and adjust the predictions. At first, the weights and biases are initialized randomly, but they are adjusted to improve the prediction as the model trains. The model undergoes training over a series of epochs, progressively adjusting its weights and biases within each epoch to minimize the errors it commits.
@@
    This resulted in a fixed set of hyperparameters (including the number of epochs) for each model architecture. These were then used to train each model on the complete dataset of 521 GRBs. This tuning process was performed separately for each model type evaluated in the study.
@@
    • Attention U-Net: Number of filters per layer, kernel size, number of convolutional blocks, learning rate, batch size, number of epochs
@@
  • Min-Max Scaling: For each GRB, the data is scaled to the [0, 1] range using standard min-max normalization (4). However, scaling is not applied to Attention U-Net and MLP models, as by performing some tuning in several GRBs showed a decrease in performance. Additionally, Attention U-Net inherently includes multiple Batch Normalization layers, while the MLP, being a relatively simpler network, is less susceptible to the weights of the model either increasing exponentially or going toward zero, namely exploding or vanishing gradients.
@@
  • 5-Fold Cross-Validation: After fixing the hyperparameters for each model, 5-fold CV is used as an alternate metric to evaluate the performance of each model. Rather than splitting the entire dataset of 521 GRBs into separate training and testing sets, we apply CV to each GRB separately. For each GRB, its time-flux data points are first randomly shuffled and then partitioned into five approximately equal-sized, non-overlapping subsets (folds). In each of the five iterations, four folds (80% of the data) are used to train the model, and the remaining fold (20%) is used to validate it (kohaviCV). This ensures that all data points are used for both training and validation across the five folds. After training, the model predicts flux values at the time points in the validation fold. These predicted fluxes are inversely transformed back to the original scale before calculating the MSE, as it depends on the scale. The MSE loss is averaged across the folds and then across the 521 GRBs. CV was used solely to evaluate model performance, not for tuning model parameters or minimizing training error.
@@ 2.3.1 Gaussian Process
GP regression (GP2003gaussian) is a probabilistic, non-parametric approach designed for regression tasks. The probabilistic nature of the GP model allows for the likelihood of the prediction instead of producing a singular prediction. The method relies on Bayesian inference, where the prior based on trends in the observed data is updated to compute a posterior that aligns with the data. The model prediction is bound by the confidence interval, which defines the region of likelihood of the prediction.
@@
In our data, out of the total sample of 521 GRBs, a tiny subset containing 9 GRBs with redshifts and 15 GRBs without redshifts, exhibited significant temporal gaps, defined as time intervals greater than 0.5 (in log10 scale). As Fig. 2a shows, the GP Regressor overfitted these data points in both the gaps, resulting in erratic predictions and extensive 95% confidence intervals. This inconsistent behavior suggested the need to fine-tune two primary hyperparameters: alpha and n_restarts_optimizer, by employing GridSearchCV; alpha is a value added to the kernel matrix’s diagonal during fitting. This hyperparameter helps avoid potential numerical issues, shown in Fig. 2a, by ensuring the matrix remains positive and definite. Furthermore, alpha can be interpreted as the additional Gaussian measurement noise variance on the training observations. The default value is 1e-10. A more considerable alpha value is frequently required to avoid overfitting in cases where the data is noisy. n_restarts_optimizer denotes the number of times the optimizer restarts to find the kernel parameters that maximize the log-marginal likelihood. This procedure guarantees a more reliable optimization of the kernel’s parameters and aids in avoiding local minima.
@@
After tuning the hyperparameters, we fit the GP Regressor to the data. The optimized model improved the reconstruction of the LCs, as shown in Fig. 2b. It might appear that Fig. 2a offers a better reconstruction than Fig. 2b, as it reconstructs the flare in the first temporal gap. However, this is misleading as very large flares like this have been unusual in the sample of the 521 GRBs. Additionally, the confidence intervals in Fig. 2a are significantly broader due to the larger temporal gap, reflecting high uncertainty. After hyperparameter optimization, the model shown in Fig. 2b not only yields smoother and more stable predictions but also exhibits reduced uncertainty, as reflected in narrower confidence intervals.
@@ 2.3.2 Multi-layer Perceptron Model
The kernel initialization of each hidden layer uses the He-Uniform (7410480) initializer method. The He-Uniform initializer initializes weights by sampling from a uniform distribution within the range:
@@
here, $n_{in}$ is the number of input units (neurons) to the layer. This ensures that the weights are small enough to prevent exploding gradients while still being large enough to avoid vanishing gradients. It is particularly suited for ReLU activation function (7410480) since it does not saturate like sigmoid or tanh activation functions. The ReLU activation function, defined as $max(0,x)$, outputs zero for negative values and keeps positive values unchanged, making it efficient for training deep networks.
@@
The model is trained using the Adam optimizer (kingma2014adam) with 0.001 as the learning rate and MSE as the loss function. The number of epochs is set to 5000. The model was implemented using the TensorFlow framework with Keras.
Next, we perform the 5-fold CV for all 521 GRBs individually and then report the average Mean Squared Error (MSE) across all GRBs as our performance metric in Table 5.
@@ 2.3.3 Attention U-Net
U-Net is a convolutional neural network architecture designed primarily for biomedical image segmentation (ronneberger2015u) but has been successfully adapted for one-dimensional data such as time series (perslev2019u). The architecture consists of a symmetric encoder-decoder structure, where the encoder compresses the input into a latent representation, and the decoder reconstructs the output from this latent space. Latent space is a lower-dimensional representation of the data where only the key features are retained and the rest discarded. Skip connections between the encoder and decoder ensure the preservation of spatial details, which are critical for high-resolution reconstruction tasks.
For the LCR, we enhance the standard U-Net with an attention mechanism(oktay2018attention), which is a section of a neural network that enhances feature extraction and boosts performance on tasks involving complicated or sequential data by dynamically focusing on the most pertinent portions of the input data and giving distinct aspects varied degrees of priority as stated in 2017arXiv170603762V. The attention mechanism enables the model to focus on relevant features at each resolution level by weighting them based on their importance. This allows the network to capture both global and local structures effectively.
@@
$\text{Attention Output}=x\cdot\psi_{f}.$ (15)
@@
The decoder reconstructs the temporal sequence by progressively restoring the temporal resolution through upsampling layers, which increases the resolution of the input data. Skip connections between the encoder and decoder layers ensure that fine-grained temporal details are preserved. To enhance the utility of skip connections, attention blocks are applied (oktay2018attention).
@@
The U-Net with Attention mechanism is well-suited for tasks involving complex temporal structures like GRB LCs, where capturing long-range dependencies and local variations is critical (perslev2019u).
@@ 2.3.4 Long Short Term Memory Neural Network
Long Short-Term Memory (LSTM; hochreiter1997long) is a form of Recurrent neural network (recurrent2001) that performs well in sequence prediction through its ability to capture long-term dependencies. Unlike traditional neural networks, LSTM incorporates feedback connections, allowing it to process data sequences rather than individual data points. This distinctive architecture equips LSTM with the ability to effectively learn and retain crucial information over time, making it a competent deep learning tool well-suited for applications in speech recognition, language translation, time series forecasting, and more (lstm2016speech; Cheng et al., 2016; lstm2019time).
@@
LSTM networks have proven valuable in astrophysics, particularly for time-series data analysis of dynamic celestial phenomena. An example is the LSTM-FCNN model, which combines LSTM and a fully connected neural network (FCNN) to measure time delays in strongly lensed Type Ia supernovae, crucial for measuring the Hubble constant (huber2024holismokes). Furthermore, LSTMs have been applied for solar flare prediction from active regions (liu2019predicting), effectively predicting them using time-series analysis. These examples show that LSTMs can outperform traditional methods in certain astrophysical time-series tasks.
Our model uses Bi-LSTM (graves2005bidirectional) to reconstruct the GRB LCs. The Bi-LSTM architecture extends the standard LSTM by processing the input data sequence in both forward and reverse directions, allowing it to capture temporal dependencies more effectively. This is particularly useful for GRB LCs, which feature complex temporal patterns such as flares and breaks. Bi-LSTMs also excel at handling missing data by learning from surrounding observations, enabling accurate imputation. The basic structure of the LSTM is described below and shown in Fig.3c. Table 2 defines all the variables and symbols.
@@ 2.3.5 Bi-Mamba Model
Mamba (gu2024mambalineartimesequencemodeling) is a novel sequence modeling architecture based on state space models (SSMs), designed to efficiently capture long-range dependencies while reducing computational complexity. Unlike long short-term memory networks (LSTMs), which rely on explicit gating mechanisms to regulate memory updates, Mamba employs a structured state space framework that enables selective long-range memory retention. This structured recurrence mechanism allows Mamba to model temporal dependencies effectively, achieving linear time complexity, making it particularly suitable for large-scale astronomical datasets such as GRB LCs, which exhibit intricate temporal variations.
@@
Since GRB LCs are discrete time-series data, the Mamba is adapted to such inputs using a Zero-Order Hold (ZOH) method – as described in (10.1093/imamci/dnac005) and (gu2024mambalineartimesequencemodeling) – which converts discrete inputs into piecewise constant signals. The resulting discrete-time formulation is given by:
@@
We used a modified version of the MambaModel from thielmann2024mambular (shown in Fig. 3d), which contains an output activation layer with LeakyReLU activation function and a linear mapping at the end. The model was built using the PyTorch framework with hyperparameter tuning using Optuna. The data (D) was scaled to a range of [0, 1] using MinMaxScaler, given in the Eq. 4. The hidden state dimension (d_state) was set to 32, while the hidden layer expansion factor (d_expand) was configured to 80. A dropout rate of 0.1 was applied to mitigate overfitting. Training was conducted over 120 epochs with a batch size of 64, using the Adam optimizer with a learning rate of 1e-4.
@@ 2.3.7 Conditional Generative Adversarial Networks
GANs goodfellow2014generative provide an innovative approach to reconstruct the LCs. Unlike traditional models that rely on Markov chains—where the current state is dependent on the previous one—GANs bypass this sequential dependency during the learning phase (mirza2014conditionalgenerativeadversarialnets), incorporating it only during gradient backpropagation. These networks are referred to as generative models and are particularly useful when dealing with complex data distributions, reducing overfitting by incorporating a dual-model approach.
@@
The CGANs mirza2014conditionalgenerativeadversarialnets extend this architecture by incorporating a condition set $y$ into the generator and discriminator, guiding the data generation process. The architecture of CGAN (as seen in Fig. 3f) comprises a generator $G(z|y)$ with two inputs: a condition set $y$ and a random noise vector $z$ and a conditioned discriminator $D(x|y)$ on $y$. The generator aims to create new samples that closely resemble the original samples by using the noise and the condition set as input. After the generation of samples, these, along with the real samples, are fed into the discriminator. The objective for the discriminator is to effectively classify between the generated and the real samples, retrospectively training the generator and discriminator in the process. It is mathematically represented as:
@@
In our implementation, the generator begins with a dense layer containing 256 neurons and the Leaky-ReLU activation function (xu2015empiricalevaluationrectifiedactivations). It includes a positive slope $m$ and comparatively a tiny negative slope $\alpha m$, preventing the occurrences of dead neurons. In our case, $\alpha$ is taken to be 1e21e-2. Following the dense layer, we utilized two Convolution-1D layers with 64 and 16 filter sizes, respectively. Paired with ReLU activation, these Convolution-1D layers enhance the feature extraction process, ensuring distinct, high-quality features. The flux and time values are preprocessed by converting them to a symmetric logarithmic scale. The time values act as condition sets for the CGAN, the flux values provide the actual data, and the noise set contains random uniform noise samples from 0 to 1 covering the whole domain. All layers in the discriminator, except the final binary classification layer, mirror those of the generator. This design ensures a balanced model complexity, preventing either network from overpowering the other. The discriminator loss is computed as the sum of binary cross-entropy losses for correctly classifying real samples as real and generated samples as fake. The generator loss is defined as the binary cross-entropy between the discriminator’s predictions on generated samples and the target label for real samples, thereby encouraging the generator to produce more realistic outputs. This results in reliable data-augmentation capabilities for the reconstruction of time series data (fu2019time; Brophy et al., 2023).
The proposed method uses TensorFlow and Keras libraries for its implementation, with 5000 epochs and a batch size of 256. The Adam optimizer was used to train the model, optimizing the loss function at a learning rate of 1e41e-4.